OpenStudy (anonymous):

find the zeros of the function h(x)=x^2+23x+60

OpenStudy (anonymous):

\[x^2+23x+60=0\] \[(x+3)(x+20)=0\] \[x+3=0\] \[x=-3\] \[x+20=0\] \[x=-20\]